# Imitation Learning-Based Online Time-Optimal Control with Multiple-Waypoint Constraints for Quadrotors

### v1

@@ II-E Polynomial-based transition phase strategy
With the proposed transition phase, for multi-waypoint flight tasks, the quadrotor can switch the waypoints flexibly without unnecessary hover so that the quadrotor can fly with aggressive maneuvers with near-time-optimal performance (Fig. 5 shows the sketch of the waypoint-switch strategy).
III Simulation result and Analysis
III-A Simulation of the WN&CNets
In this simulation, we utilize the dynamic model from Equation 2 as the quadrotors’ model. The inputs for the simulation are the total thrust command $T_{c}$ and the torque command $\boldsymbol{\tau}_{c}$. The WN&CNets provides the total thrust command $T_{c}$ and the angular rate command $\boldsymbol{\omega}_{c}$. We use a proportional control (P control) to calculate the torque commands $\boldsymbol{\tau}_{c}$ and send them to the dynamic model.
To illustrate the time optimality of the WN&CNets. We first compare the WN&CNets to the CPC method in a simple two-waypoint task, where the start point is $(0.0,0.0,1.0)$. The waypoints is $(3.6,1.2,1.4)$ and the endpoint is $(6.9,0.0,1.0)$ (Fig. 6). Fig. 7 illustrates the position curves generated by the CPC method and the WN&CNets, as well as the rate commands provided by the CPC and the WN&CNets.
Figure 6: 3D trajectories generated by the WN&CNets and the CPC in simulation.
Figure 7: Comparison of the proposed method and the CPC method for a two-waypoint flight task.
From the simulation results, it can be seen that for the same flight task, the trajectories of the WN&CNets are very close to the one generated by the CPC method (the difference between arrival time is $0.01s$). Although the quadrotor using the WN&CNets arrives at the final waypoint with a $10cm$ error in position compared to the CPC method, it achieves online performance in return, which also indicates that the proposed method achieves the online near-time-optimal performance.
III-B Polynomial-based transition phase strategy and comparison
In this section, we design two experiments to verify the performance of the transition strategies for the two scenarios proposed in section E. The position of the waypoints are shown in TABLE I.
TABLE I: The position of the waypoints in the simulation experiment
Experiment 1 Experiment 2
Start point $(0.0,0.0,1.0)$ $(-3.0,0.0,1.0)$
Waypoint 1 $(3.5,-1.0,1.5)$ $(0.0,1.0,1.5)$
Waypoint 2 $(6.5,0.0,1.0)$ $(3.0,0.0,1.0)$
Waypoint 3 $(10.0,1.0,0.6)$ $(0.0,-1.0,0.5)$
Waypoint 4 $(14.0,2.0,1.0)$ $(-3.0,0.0,1.0)$
In Experiment 1, we set $v_{in}=100\%$, $90\%$, $80\%$$v_{max}$ and in Experiment 2, $a_{in}$ are set to $100\%$, $90\%$, and $50\%$$a_{max}$ to test the performance of different entering conditions, where $v_{max}$ and $a_{max}$ are the maximum velocity and acceleration of the quadrotor respectively. Also, we use the CPC method and the classic Minimum Snap Trajectory Generation and Control method (MSTG&C) [4] as the benchmark to compare the performance of the proposed method. The results are shown in Fig. 8 and the arrival time of each method is listed in Table II. To demonstrate the difference of arrival time $t_{f}^{*}$ among each method, we introduce an index $\sigma=(t_{f}^{*}-t_{f}^{cpc})/t_{f}^{cpc}$. The higher $\sigma$ is, the slower the quadrotor flies compared to the CPC method.
(a) Experiment 1: same direction flight
(b) Experiment 2: opposite direction flight
Figure 8: Comparison of the proposed method, the CPC method and the MSTG&C method in multiple waypoint tasks.
TABLE II: Arrival Time of each method
Experiment 1 Experiment 2
  WN&CNet $v_{in}$ $t_{f}$[s] $\sigma$[%] $a_{in}$ $t_{f}$[s] $\sigma$[%]
$stop~{}and~{}go$ 4.14 62.4 $stop~{}and~{}go$ 4.15 30.9
$100\%v_{max}$ 2.80 9.8 $100\%a_{max}$ 3.46 9.15
$90\%v_{max}$ 2.87 12.5 $90\%a_{max}$ 3.58 12.9
$80\%v_{max}$ 2.98 16.9 $50\%a_{max}$ 3.81 20.2
  MSTG&C / 3.25 27.5 / 4.13 30.3
  CPC / 2.55 0 / 3.17 0
From the simulation, we can find that firstly, compared to the origin WN&CNets without the transition phase, the WN&CNets with transition can significantly reduce the flight time. It is also clear that with fine-tuned parameters, the flight time of the WN&CNets can surpass the MSTG&C while the latter needs offline optimization to achieve near-time-optimal performance. Compared to the CPC approach, the best result of the proposed method is slightly slower (less than 10% in both Experiment 1 and Experiment 2). However, it should be noted that the CPC method usually needs dozens of minutes or more to calculate one time-optimal trajectory while the WN&CNets can achieve online performance, which makes it more suitable for practical and flexible real-world applications.
Indeed, expanding the dataset of optimal trajectories by incorporating more waypoints or a wider range of start and end states in the training data can enhance the performance of the network controller. However, this expansion process can be costly and may never reach a finite point of completeness. To strike a balance between optimality and effectiveness, it is beneficial to incorporate an additional planning and control strategy, such as polynomials in this work, to handle situations where the quadrotor ventures beyond the safe state space covered by the training data.
IV Experiment setup and result
IV-A Experiment setup
@@
For the real-world experiments, we utilize a self-made quadrotor as the flying platform. This quadrotor weighs only $100$ $g$ and boasts a thrust-to-weight ratio of about $3$, as shown in Fig. 9. To have precise and stable angular velocity tracking, the quadrotor is equipped with a $10.0$ $g$ autopilot with the Betaflight open-source software running onboard, which serves as an angular rate loop controller. The WN&CNets, running on a laptop, feed angular velocity commands and throttle commands to the quadrotor by the MAVLink protocol via an onboard WiFi module.
Figure 9: The quadrotor vehicle
To acquire accurate measurements of the quadrotor’s states, the onboard autopilot transmits IMU measurements to the laptop via the onboard WiFi module. Additionally, the Optitrack motion capture system is used to provide accurate position measurements of the quadrotor. On the laptop, the quadrotor’s position, velocity, acceleration, and angular velocity are sent to the WN&CNets to predict the angular velocity and thrust commands. These commands are sent back to the autopilot via the WiFi module to steer the quadrotor to fly toward the next waypoints.
In our real-world experiments, to ensure safety and allow for a necessary margin, the free-flying space for the quadrotor to maneuver is $6.0$ $m\times 4.0$ $m\times 2.0$ $m$. Due to the limited flying space, the quadrotor is programmed to execute flights with $7$ waypoints to demonstrate the performance of the proposed method in real-world scenarios.
IV-B Experiment results
(a) The control structure of the WN&CNets
(b) The control structure of the transition phase and the CPC
Figure 10: The control structure of the WN&CNets, the transition phase and the CPC.
Figure 11: The flying experiment of the quadrotor flying through moving waypoints.
TABLE III: The position of the waypoints and flight results
Flight ID Start Point Waypoint Endpoint Flight Time [s] Error [m]
@@ IV-B Experiment results
$1$ $-3.0$ $-1.0$ $1.0$ $1.5\cos\frac{\pi}{8}t$ $1.01.5\sin\frac{\pi}{8}t-$ 1.0 $0.6$ $3.0$ $1.0$ $1.0$ $2.65$ $0.36$
$2$ $3.0$ $1.0$ $1.0$ $1.5\cos\frac{\pi}{8}t$ $1.01.5\sin\frac{\pi}{8}t+$ 1.0 $1.4$ $-3.0$ $1.0$ $1.0$ $2.57$ $0.19$
$3$ $-3.0$ $1.0$ $1.0$ $1.5\cos\frac{\pi}{8}t$ $1.01.5\sin\frac{\pi}{8}t+$ 1.0 $1.4$ $3.0$ $-1.0$ $1.0$ $2.34$ $0.07$
$4$ $3.0$ $-1.0$ $1.0$ $1.5\cos\frac{\pi}{8}t$ $1.01.5\sin\frac{\pi}{8}t-$ 1.0 $0.6$ $-3.0$ $-1.0$ $1.0$ $2.45$ $0.16$
The control loop is shown in Fig. 10 (a) where the WN&CNets runs on a laptop and is fed with the measured quadrotor’s states. The predicted commands are sent back to the quadrotor to do the maneuver. As for the transition phase and the CPC, a classic differential flatness-based feed-forward and feedback strategy is employed, which is shown in Fig. 10 (b).
Figure 12: Trajectories of the quadrotor flying through moving waypoints via WN&CNets.
The first experiment is designed to let the quadrotor fly through moving waypoints to demonstrate that the proposed method can handle dynamic waypoints online. Table III lists the start points’ position, moving waypoints and endpoints. In this experiment, the quadrotor will first hover at the start points and then use the WN&CNets to fly through the waypoints in a pre-defined sequence and finally hover at the endpoints. During the flight, the quadrotor only knows the current position of the waypoints instead of their whole trajectories. The current positions of the endpoints and moving waypoints are fed to the WN&CNets at the frequency of $30$ $Hz$. As a result, the WN&CNets can steer the quadrotor to fly through the moving waypoints and hover at the endpoints with its aggressive maneuverability. The flight time and the errors between the quadrotor and the moving waypoints are also listed in Table III. The experiment scenarios are presented in Fig. 11 where we use blue and red gates to represent the moving waypoints. The trajectories of the quadrotor together with the moving waypoints are depicted in Fig. 12. From the experiments, it can be seen that by only knowing the current position of the waypoints, the quadrotor can fly through these moving waypoints and hover at the endpoints. Thus, compared to the CPC, the proposed method can handle the moving waypoints online while maintaining its time optimality.
We also design an experiment to test the proposed transition phase strategy in the real world, where the quadrotor first hovers at the start point, then flies through $7$ waypoints aggressively and finally hovers at the endpoint. The position of each waypoint is listed in Table IV.
TABLE IV: The position of the waypoints in the real-world experiment
@@
$(3.0,0.0,1.0)$ $(0.0,-1.0,0.6)$ $(-3.0,-2.0,1.0)$
Figure 13: Trajectories of the quadrotor flying through $7$ waypoints via WN&CNets and transition phase strategy.
At the same time, we use the CPC method to steer the quadrotor to fly through the same track as the benchmark. In the experiments, both the CPC and our WN&CNets can control the quadrotor to fly through all the waypoints successfully and the quadrotor can achieve a maximum speed of $7.0$ $m/s$. To demonstrate the difference in the experiment results between the two methods, we list the arrival time $t_{f}$, the maximum velocity $v_{max}$, the maximum error $E_{max}$ when passing through waypoints, and the processing time for planning $T_{p}$ in Table V. The experiment scenarios are depicted in Fig. 1 and the trajectories of the quadrotors are shown in Fig. 13.
TABLE V: Experiment results: comparison between each method
@@
From the real-world experiments, we can see that although the arrival time of the proposed method ($7.18$ $s$) is slightly larger than the CPC method ($6.85$ $s$), the processing time of the proposed method and the CPC method differs by orders of magnitude, where the CPC method needs $11min$ to calculate the time-optimal trajectories while the proposed method can be run at the frequency of $30$ $Hz$ (the frequency can actually be much higher, but this is the frequency that is used in our experiments) to predict the time-optimal trajectories. In terms of the tracking error, due to the presence of unknown disturbances and model inaccuracy, the CPC trajectory, which can not be run online, passes through waypoints with larger errors ($0.47$ $m$) than the proposed method ($0.26$ $m$). To conclude, compared to the CPC method, in multi-waypoint flight, although the proposed method slightly loses time optimality, it significantly surpasses the CPC method in its efficiency which makes it more practical and flexible for most flight tasks.
V CONCLUSIONS
@@
In this letter, we develop a novel network WN&CNet which can steer the quadrotor to fly through multiple waypoints with the minimum time. Due to the quadrotors’ high-dimension state space and very limited time-optimal training data, the trained neural network can only guide the quadrotor fly through $2$ waypoints and hover at the second waypoint. To address this issue, we develop a transition phase strategy to help the quadrotor jump over the so-called stop-and-go phase. So that the quadrotor can consecutively fly through multiple waypoints with its aggressive maneuverability. We also conduct simulations and real-world experiments to test our method. The results indicate that although the proposed method slightly loses the time-optimality compared to the CPC method, it gains the ability to be run online with $30Hz$ in return, which makes it more practical for real-world flying tasks that need aggressive maneuvers.
@@
There are several directions for future work. The real inputs of the quadrotors, the rotors’ speed, can be added to the quadrotor’s model and the optimization problem to further improve the time optimality. The sensors’ delay would be considered in our future work so that they can be estimated and compensated online to eliminate the negative effects on the network’s prediction accuracy. Two-drone racing with the proposed WN&CNets can also be an interesting topic to research on.

### v2

@@ II-E MINCO-based transition phase strategy
When the boundaries of the MINCO trajectories are determined (positions, velocities, and accelerations), the transition phase can be calculated. With the proposed transition phase, for multi-waypoint flight tasks, the quadrotor can switch the waypoints flexibly without unnecessary hover so that the quadrotor can fly with aggressive maneuvers with near-time-optimal performance (Fig. 4 shows the sketch of the waypoint-switch strategy).
III Experiment Result and Analysis
In this section, we will analyze the time optimality and real-time capability of the proposed WN&CNets through both simulation and real-world experiments. For simulation experiments, we will introduce CPC [11] as the benchmark and compare a great amount of CPC trajectories with the WN&CNets for executing the same waypoint-constrained flight. For real-world experiments, we first use the WN&CNets to navigate the quadrotor in two-waypoint flights with the first waypoint fixed or moving continuously. Next, we will demonstrate the performance of the WN&CNets and MINCO transition through various multi-waypoint experiments.
III-A Experiment setup
Figure 5: The quadrotor vehicle.
@@ III-A Experiment setup
For real-world experiments, we utilize a self-made quadrotor as the flying platform, which weighs less than $200$g and boasts a thrust-to-weight ratio of about $2.2$, as shown in Fig. 5. To have precise and stable angular velocity tracking, the quadrotor is equipped with a $10$g autopilot with the Betaflight open-source software running onboard, which serves as an angular rate loop controller. The quadrotor is equipped with a Cool Pi computer with the proposed WN&CNets running onboard. The rate commands and throttle commands are fed to the autopilot using the MAVLink protocol via the serial port. In our real-world experiments, to ensure safety and allow for a necessary margin, the free-flying space for the quadrotor to maneuver is set to be $5.5m\times 5.5m\times 2.0m$.
To acquire accurate measurements of the quadrotor’s states, the Optitrack motion capture system is used to provide accurate position and attitude measurements. To evaluate the delay of the measurements, we sum the delay caused by the motion capture system’s calculation time which is provided by the Optitrack software and the delay of transmitting the data from the Optitrack system to the Cool Pi. The total delay is about $30ms$, which is considered in our network training.
On the Cool Pi, the WN&CNet is deployed with LibTorch, and can achieve a maximum prediction frequency of $950$Hz. In our experiments, the quadrotor’s position, velocity, attitude, and total thrust are sent to the WN&CNets to predict the rate and throttle commands at $100$Hz. The predicted commands are sent to the autopilot to perform the maneuver.
As for the transition phase (MINCO) and the benchmark (CPC), a classic differential flatness-based feed-forward and feedback strategy [8] is employed.
III-B Time optimality of the WN&CNets
Figure 6: Part of 3D trajectories generated by the CPC and the WN&CNets in simulation.
Figure 7: The navigation results of the WN&CNets in simulations with different delay.
To illustrate the time optimality of the WN&CNets, we first compare the WN&CNets with the CPC method in 1000 simple two-waypoint tasks distributed in the space of $[-5m,5m]\times[-5m,5m]\times[0.4m,2m]$. Part of the 3D trajectories generated by the CPC and the WN&CNet are shown in Fig. 6. In the simulation, we utilize the dynamic model from Equation 2 as the quadrotors’ model. The time step of the simulation is $0.025s$. The inputs for the simulation are the total thrust command $T_{c}$ and the torque command $\boldsymbol{\tau}_{c}$. The WN&CNet provides the total thrust command $T_{c}$ and the angular rate command $\boldsymbol{\omega}_{c}$. We use a proportional control to calculate the torque commands $\boldsymbol{\tau}_{c}$ and send them to the dynamic model. As for the CPC, we directly use the trajectories generated by the CPC method.
We use a tolerance of $0.2m$ to determine if the quadrotor passes the waypoint. The simulation result shows that in these $1000$ flight tasks, the average time of the CPC trajectories is $1.89s$, while the average time for the quadrotor to complete the tasks navigated by the WN&CNets is $1.91s$ (just one simulation step longer than CPC). The average minimum error in passing through the first waypoint by the CPC is $0.19m$ (just within the tolerance range), while by the WN&CNets is $0.24m$.
To illustrate the influence of the delay on flight performance, we introduce different delays in one of the aforementioned flights. The navigation results of the WN&CNets with the ideal CPC are shown in Fig. 7. The quadrotor achieves the closest arrival time to the CPC in the simulation with a delay of $30ms$, which is the delay considered in the network training. It can also be found that with the larger delay mismatch, it is challenging for the same WN&CNets to navigate the quadrotor.
Figure 8: Comparison of the proposed method and the CPC method for a two-waypoint flight task in real-world experiments.
TABLE I: The position of the waypoints and flight results
Flight Start Point Waypoint Endpoint Maximum Error [m]
@@ III-B Time optimality of the WN&CNets
$1$ $-2.5$ $1.0$ $1.0$ $0.5\sin\frac{\pi}{5}t$ $1.50.5\cos\frac{\pi}{5}t+$ 1.5 $1.5$ $2.5$ $1.0$ $1.0$ $0.39$
$2$ $2.5$ $1.0$ $1.0$ $0.5\sin(\frac{\pi}{5}t+1.0)$ $0.5\cos(\frac{\pi}{5}t+1.0)$ $1.0$ $-2.5$ $-1.0$ $1.0$ $0.29$
$3$ $-2.5$ $-1.0$ $1.0$ $0.5\sin(\frac{\pi}{5}t+1.5)$ $1.50.5\cos(\frac{\pi}{5}t+1.5)-$ 1.5 $0.5$ $2.5$ $-1.0$ $1.0$ $0.37$
$4$ $2.5$ $-1.0$ $1.0$ $0.5\sin(\frac{\pi}{5}t+1.0)$ $0.5\cos(\frac{\pi}{5}t+1.0)$ $1.0$ $-2.5$ $1.0$ $1.0$ $0.25$
(a) The scenarios of the quadrotor passing through the moving waypoints for 4 flights. We use blue, green and red gates to represent the moving waypoints.
(b) The trajectories and velocity profiles of the quadrotor maneuvering for $3$ laps.
Figure 9: The scenarios and the trajectories of the quadrotor flying through moving waypoints navigated using the WN&CNets.
We also perform the comparison of the WN&CNets and CPC in real-world flights. Fig. 8 demonstrates the position curves generated by the CPC method and the WN&CNets in one specific two-waypoint task, in which the start point is $(-2.55,1.0,1.0)$ and the waypoints are first $(0.0,0.25,1.4)$ and then $(2.5,1.1,1.0)$. The result indicates that for the same flight task, the trajectories of the WN&CNet are very close to the one generated by the CPC method, while achieving real-time computing. It indicates that the proposed method achieves the online near-time-optimal performance.
The second real-world experiment is designed to let the quadrotor fly through moving waypoints to demonstrate that the proposed method can handle dynamic waypoints online. Table I lists the positions of the start points, moving waypoints and endpoints. In this experiment, the quadrotor will first hover at the start points and then use the WN&CNets to fly through the continuously moving waypoints and finally hover at the endpoints. During the flight, the quadrotor only knows the current position of the waypoints instead of their whole trajectories. The positions of the waypoints are fed to the WN&CNets at the frequency of $100$Hz. In this experiment, the quadrotor maneuvers in this manner for $3$ laps. The maximum errors of all the laps between the quadrotor and the moving waypoints are also listed in Table I. The experiment scenarios and trajectories are presented in Fig. 9. This experiment indicates one important advantage over the benchmark, CPC, that is due to the neural networks’ high efficiency and high-frequency prediction, the proposed method is able to handle the dynamic waypoints online while the CPC method needs several seconds to several minutes to optimize a feasible trajectory which makes it impossible to run online and handle dynamic waypoints during the flight.
III-C WN&CNets and transition strategy for multi-waypoint flight tasks
In this section, we design an experiment to test the proposed transition strategy. In this experiment, the quadrotor first hovers at the start point, then flies through $7$ waypoints aggressively and finally hovers at the endpoint (Fig. 10). The positions of each waypoint are listed in Table II.
TABLE II: Waypoints’ position for multi-waypoint experiment
@@ III-C WN&CNets and transition strategy for multi-waypoint flight tasks
$(2.2,0.4,1.0)$ $(0.0,-0.4,0.6)$ $(-2.2,-1.5,1.0)$
Figure 10: Trajectories of the quadrotor flying through $7$ waypoints via WN&CNets and transition phase strategy along with the benchmarks. The red numbers represent the order of the waypoints.
At the same time, we use the CPC method and the MINCO method to steer the quadrotor to fly the same track as the benchmarks. In the experiments, the CPC, the MINCO and our WN&CNets can navigate the quadrotor to fly through all the waypoints successfully and the quadrotor can achieve a maximum speed of up to $5.0m/s$. The trajectory curves generated by the three methods are shown in Fig. 11. To compare the performance of the three methods, we list the arrival time $t_{f}$, the maximum velocity $v_{max}$, the maximum error $E_{max}$ of passing through the waypoints, and the processing time $T_{p}$ of generating the optimal control commands in Table III.
From Table 3, it can be seen that although the arrival time of the proposed method ($7.09s$) is slightly longer than the CPC method ($6.66s$), the processing time of the proposed method and the CPC method differs by orders of magnitude. The CPC method needs $9min$ to calculate the time-optimal trajectories while the proposed method can be run at the frequency of $100$Hz (the WN&CNets can achieve a prediction frequency of $950$Hz, but $100$Hz is the frequency that is used in our experiments) to predict the time-optimal trajectories. In terms of the tracking error, due to the tolerance of $0.2m$ in the optimization problem as well as the presence of unknown disturbances and model inaccuracy, the CPC trajectory, which can not replan online, passes through waypoints with maximum error as $0.29m$.
Figure 11: Comparison of the proposed method, CPC and MINCO for a multi-waypoint flight task in real-world experiments.
TABLE III: Experiment results: comparison between each method
@@
The MINCO can be solved online and analytically passes through each waypoint accurately, which means that the errors are completely caused by trajectory tracking errors. The maximum error of the WN&CNets along with the MINCO transition is $0.22m$, which is slightly larger than the tolerance for generating the time optimal datasets while it is close to the average error in the simulation. To conclude, in multi-waypoint flight tasks, although the proposed method slightly loses time optimality, it significantly surpasses the CPC method in its efficiency which makes it more practical and flexible.
To illustrate the feasibility of the proposed method to be applied to a wider range of scenarios, we also demonstrate that the quadrotor can fly with flower-like trajectories (Fig. 12) and continuously pass through moving target points without hovering (Fig. 1).
Figure 12: Flower-like trajectories generated by the proposed method (a lap of 11-petal trajectory contains 22 waypoints).
III-D WN&CNets with headings
As we discussed in Section II-B, for some flight tasks, the heading needs to be specified at the waypoints. Although in this letter, we mainly focus on time-optimal maneuvers with free heading, we also generate the dataset using constraint (5) and train another WN&CNets with heading constraints. In terms of training, the input heading measurement of the network should be changed into a relative heading to the desired one, which is similar to the relative position inputs. The flying results show that the WN&CNets with headings can steer the quadrotor pass through the waypoints with specified headings in the minimum time. The pictures and the yaw curves of the experiments are shown in Fig. 13.
Figure 13: The quadrotor is steered by the WN&CNets (compared with the CPC) to pass through the first waypoint with a heading of $-60^{\circ}$ and hover at the final waypoint with a heading of $-0^{\circ}$ (the yellow arrow indicates the heading of the quadrotor).
IV CONCLUSIONS
@@
In this letter, we develop a novel network WN&CNet which can navigate the quadrotor to fly through multiple waypoints with the minimum time. Due to the quadrotors’ high-dimension state space and very limited time-optimal training data, the trained neural network can only guide the quadrotor fly through $2$ waypoints and hover at the second waypoint. To address this issue, with the aid of MINCO trajectories, we develop a transition phase strategy to help the quadrotor jump over the so-called stop-and-go phase. So that the quadrotor can consecutively fly through multiple waypoints with its aggressive maneuverability. We also conduct simulations and real-world experiments to test our method. The results indicate that although the proposed method slightly loses the time-optimality compared to the CPC method, it gains the ability to be run online with $100$Hz in return, which makes it more practical for real-world flying tasks that need aggressive maneuvers.
@@
There are several directions for future work. The real inputs of the quadrotors, the rotors’ speed, can be added to the quadrotor’s model and the optimization problem to further improve the time optimality. The sensors’ delay would be considered in our future work so that they can be estimated and compensated online to eliminate the negative effects on the network’s prediction accuracy. More computationally efficient time-optimal trajectory generation methods can be adopted to replace the CPC to generate datasets, so that we can acquire much more training data in confined time and expand the scope of application of the WN&CNets. Two-drone racing with the proposed WN&CNets can also be an interesting topic to research on.